OpenStudy (anonymous):

A 4.0kg box slides a 20m distance on ice. If the coefficient of kinetic friction is 0.22, what is the work done by the friction force?

OpenStudy (anonymous):

\[ W=F_f\cdot d\cdot\cos\theta\\ F_f=\mu N\quad,\quad d=20\quad,\quad\theta=180^\circ \]